A turtle crawls along a straight line, which we will call the 𝑥-axis with a positive direction to the right. The equation for the turtle’s position as a function of time is 𝑥(𝑡) = 50.0𝑐𝑚 + (12.00𝑐𝑚/𝑠 2 )𝑡. (a) Find the turtle’s initial velocity, initial position, and initial acceleration. (b) At what time 𝑡 is the velocity of the turtle zero?


1
Expert's answer
2021-09-26T18:47:10-0400

Given:

"x(t)=50.0+12.00t^2\:\rm (cm)"


The turtle’s velocity

"v(t)=x'(t)=24.00t\:\rm cm/s"


The turtle’s acceleration

"a(t)=v'(t)=24.00\:\rm cm/s^2"

(a) The turtle’s

initial velocity:

"v(0)=0\:\rm cm/s"

 initial position:

"x(0)=50.0+12.00*0^2=50\:\rm cm"

initial acceleration:

"a(0)=24.00\:\rm cm/s^2"


(b) The velocity of the turtle is zero at the instant

"t=0\:\rm s"